# Quantum Key Leasing for PKE and FHE with a Classical Lessor

### 03 Jan 2024 v3

@@ 3 Acknowledgement
We thank an anonymous reviewer who points out an issue in the analysis of the LWE search-to-decision extractor analysis, which has now been fixed.
@@ 5.1 Projective Implementation and Threshold Implementation
Motivated by the discussion above, [Zha20] (inspired by [MW05]) formalizes a new measurement procedure for testing a state received by an adversary. We will be adopting this procedure when defining security of secure key leasing schemes in Section Section 8. Consider the following procedure as a binary POVM ${\mathcal{P}}$ acting on an alleged quantum-decryptor program $\rho$: sample a ciphertext $x$, evaluates the quantum decryptor on $x$, and checks if the output is correct. In a nutshell, the new procedure consists of applying an appropriate projective measurement which measures the success probability of the tested state $\rho$ under ${\mathcal{P}}$, and to output “accept” if the success probability is high enough. Of course, such measurement will not be able extract the exact success probability of $\rho$, as this is impossible from we have argued in the discussion above. Rather, the measurement will output a success probability from a finite set, such that the expected value of the output matches the true success probability of $\rho$. We will now describe this procedure in more detail.
@@ 5.2 Approximating Threshold Implementation
Our algorithm is parameterized by a distribution ${\mathcal{D}}$, collection of projective measurements ${\mathcal{P}}$, and real values $0<\epsilon,\delta,\gamma\leq 1$, and is denoted as ${\sf ATI}_{{\mathcal{P}},{\mathcal{D}},\gamma}^{\epsilon,\delta}$. The running time of the following algorithm is $\frac{2\ln(4/\delta)}{\epsilon^{2}}$.
@@
Unfortunately these above two theorems do not possess precise enough error parameters for our use in some settings. We additionally show a theorem with more precise parameters for two statistically close distributions.
Theorem 5.11.
let ${\mathcal{D}}_{0},{\mathcal{D}}_{1}$ be two statistically-close distributions over ${\mathcal{R}}$, with distance $\eta$. Then for any inverse polynomial $\epsilon$ and any function $\delta$, it holds that:
$\operatorname{Tr}[{\sf ATI}_{{\mathcal{D}}_{1},{\mathcal{P}},\gamma}^{\epsilon,\delta}(\rho)]\geq\operatorname{Tr}[{\sf ATI}_{{\mathcal{D}}_{0},{\mathcal{P}},\gamma}^{\epsilon,\delta}(\rho)]-O({\sf poly}(\lambda)\cdot(\log(1/\delta)/\epsilon^{2})^{3})\cdot(\frac{1}{|{\mathcal{R}}|}+\eta^{\frac{3}{2}})$
Since we work with $|{\mathcal{R}}|$ with exponential size, we will assume $\frac{1}{|{\mathcal{R}}|}$ to be suppressed when we use the theoreom.
We refer the proof or the above theorem to Section C.2
@@ 5.2.1 Identity-like Approximate Threshold Implementation
  • For any input state $\rho$, let $\rho^{\prime}$ be the state after applying ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},1/2+\gamma}$ on $\rho$, we have $\lVert\rho-\rho^{\prime}\rVert_{\operatorname{Tr}}\leq O(\eta\cdot\frac{\ln(4/\delta)}{\epsilon})$  [^7] In our applications, $\eta\cdot\frac{\ln(4/\delta)}{\epsilon}$ is smaller than inverse exponential in security parameter $\lambda$.
@@
We obtain this theorem by making observations on the state after running the ATI algorithm on page 20 of our submission.
We will use Claim 6.3 in [Zha20], and some explanations under theorem 6.2 in [Zha20].
We can write the input state at the beginning as $\rho=\sum_{i}\lambda_{i}\ket{\phi_{i}}\bra{\phi_{i}}$ |, where $p_{i}$ is an eigenvalue with eigenvector |$\ket{\phi_{i}}$ for measurement ${\mathcal{P}}_{{\mathcal{D}}_{0},{\mathcal{D}}_{1}}$ in Definition 5.12(see discussions inSection 5.1 and Lemma 5.2). In other words, ${\mathcal{P}}_{{\mathcal{D}}_{0},{\mathcal{D}}_{1}}\ket{\phi_{i}}=p_{i}\ket{\phi_{i}}$. Meanwhile, since {|$\{\ket{\phi_{i}}\}_{i}$ form a basis, {|$\{\ket{\phi_{i}}\}_{i}$ is also a basis for the density operator $\rho$ with eigenvalues $\{\lambda_{i}\}_{i}$ because such a spectral decomposition always exists for any $\rho$.
@@
We consider the state after doing $T$ number of iterations of the main loop in ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},1/2+\gamma}$ algorithm. Let $L=(L_{1},L_{2},\cdots,L_{2T})$ be the sequence of measurement outcomes we obtain in the ATI algorithm from the alternating tests ${\sf CProj}_{{\mathcal{P}}}$ and ${\sf IsUniform}$; let $t$ be the number of $i$ such that $L_{i-1}=L_{i}$ we have in $L$.
Since at the end of the algorithm ${\sf ATI}$, we will stop at a ${\sf IsUniform}$ test with measurement outcome 1, we obtain a state following [Zha20] Claim 6.3 as follows.
$|\displaystyle\rho^{\prime}=\sum_{i}\lambda_{i}\cdot\frac{p_{i}^{t}\cdot(1-p_{i})^{2T-t}}{\sum_{i}\lambda_{i}\cdot\cdot p_{i}^{t}\cdot(1-p_{i})^{2T-t}}\ket{\phi_{i}}\bra{\phi_{i}}$ | (3)
[Zha20] gave the form of an unnormalized pure state w.l.o.g.; we can easily obtain that the form of a (normalized) mixed state will be the above.
@@
Here $\tilde{p}=\frac{t}{2T}$ is the value we obtain at the next-to-last step in ATI algorithm in Section 5.2 (before we check $\tilde{p}\geq 1/2+\gamma$); $p$ is an outcome we obtain if we apply an ideal projective implementation of ${\mathcal{P}}$– in short we obtain some $p=p_{i^{*}}$ with probability $\sum_{j}\lambda_{j}$ where all |$\ket{\phi_{j}}$ have eigenvalues $p_{j}=p_{i^{*}}$ under ${\mathcal{P}}$ (see discussions in Section 5.1). But we don’t need to worry about the detailed distribution of $p$ here: since we must have $p\in[1/2,1/2+\eta]$, we can deduce that $\frac{(1-\epsilon)}{2}\cdot 2T\leq t\leq(\frac{(1+\epsilon)}{2}+\eta)\cdot 2T$. Since we have $\eta<\epsilon/2$ by our assumption and let $T=\ln(4/\delta)/\epsilon^{2}$ we have $T-\frac{\ln(4/\delta)}{\epsilon}\leq t\leq T+\frac{2\ln(4/\delta)}{\epsilon}$, with probability $1-\delta$. Now we analyze the trace distance between $\rho,\rho^{\prime}$. Without loss of generality, first assume $t\geq T$ and $\lambda_{i}-\lambda_{i}\cdot\frac{p_{i}^{t}\cdot(1-p_{i})^{2T-t}}{\sum_{i}\lambda_{i}\cdot p_{i}^{t}\cdot(1-p_{i})^{2T-t}}\geq 0,\forall i$.
@@
$\leq\frac{1}{2}\sum_{i}\lambda_{i}(\cdot 1-(1-4\frac{\ln(4\delta)}{\epsilon^{2}}\cdot\eta^{2}-8\eta\cdot\frac{2\ln(4\delta)}{\epsilon}-4\eta^{2}\cdot 8\eta\cdot\frac{2\ln(4\delta)}{\epsilon}))\text{ since }T=\frac{\ln(4/\delta)}{\epsilon^{2}}=(T-t)/2\epsilon\allowdisplaybreaks$
@@
$\lVert\rho-\rho^{\prime}\rVert_{\operatorname{Tr}}$ $\leq\frac{1}{2}\sum_{i}\lambda_{i}((1-4\eta^{2})^{T}\cdot(\frac{1/2+\eta}{1/2-\eta})^{2(t-T)}-1)$
$\leq\frac{1}{2}\sum_{i}\lambda_{i}((1-4T\eta^{2})\cdot(1+4\eta(2t-2T))-1)$
@@
Suppose the statistical distance of ${\mathcal{D}}_{0},{\mathcal{D}}_{1}$ is 0. ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},1/2+\gamma}$ is the ATI for mixture of projections ${\mathcal{P}}=({\mathcal{P}}_{{\mathcal{D}}_{0},{\mathcal{D}}_{1}},{\mathcal{Q}}_{{\mathcal{D}}_{0},{\mathcal{D}}_{1}})$ defined in Definition 5.12, with parameters $\gamma,\epsilon,\delta$ where $\gamma$ is inverse polynomial and $\epsilon<\gamma$. For any input state $\rho$, let $\rho^{\prime}$ be the state after applying ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},1/2+\gamma}$ on $\rho$, we have that $\lVert\rho-\rho^{\prime}\rVert_{\operatorname{Tr}}=0$.
@@ 7.2.4 Efficient Range Preparation for NTCF with Small LWE Secrets
Next, we measure the last register to obtain $\mathbf{y}=\mathbf{x}_{b}\mathbf{A}+\mathbf{e}^{\prime}+\mathbf{b}\cdot\mathbf{s}\mathbf{A}$. We can then uncompute the register containing $\mathbf{e}^{\prime}$ using the information in register containing $\mathbf{x}$, the key $(\mathbf{A},\mathbf{u})$ and the last register.
@@ 8.2 Strong SKL-PKE PKE Security: Threshold Implementation Version
In this section, we define a security notion we call Strong SKL-PKE, which is described via the measurement ${\sf TI}$ in Section 5. We show that it implies the regular security notion in the previous section.
@@ 9.1 Parameters
    The ratios $\frac{B_{V}}{B_{L}}$, $\frac{B_{P}}{B_{V}}$, $\frac{B_{P^{\prime}}}{B_{P}}$, $\frac{B_{X}}{B_{S}}$, and $\frac{B_{P^{\prime}}}{B_{X}^{4}\cdot B_{P}}$ are super-polynomial in $\lambda$.
@@ 9.3 Correctness
Note that we can write out the entire quantum key as ${\sf\rho_{\mathsf{sk}}}=\frac{1}{\sqrt{2^{k}}}\sum_{\mathbf{b}_{j}\in\{0,1\}^{k}}\lvert\mathbf{b}_{j,1},\mathbf{x}_{k,\mathbf{b}_{j,1}},\cdots,\mathbf{b}_{j,k},\mathbf{x}_{1,\mathbf{b}_{j,k}}\rangle$ By applying the decryption procedure, we will have:
@@
Since we have $\lVert\sum_{i}(\mathbf{e}_{i}^{\prime}+\mathbf{e}_{i})\rVert\leq 2k\cdot\sqrt{m}\cdot B_{P},\lVert\mathbf{e}^{\prime\prime}\rVert\leq\sqrt{m}\cdot B_{P^{\prime}}$ and $\lVert(\sum_{i\in[k]}(-\mathbf{b}_{j,i}\cdot\mathbf{e}_{i}+\mathbf{e}_{i}^{\prime})\cdot\mathbf{R}+\mathbf{e}^{\prime\prime})\mathbf{G}^{-1}(\mathbf{0}|\lfloor\frac{q}{2}\rfloor)\rVert\leq\frac{q}{4}$ for all support $\mathbf{b}_{j}\in\{0,1\}^{k}$ in the final outcome, we will obtain $\mu$ with all but negligible probability, for all support in the above state. We can write the final output for $\mu$ in the third (rightmost) register and use $\mathsf{ct}$ to uncompute the second register and recover ${\sf\rho_{\mathsf{sk}}}$.
@@
Reusability of the decryption key follows from correctness and the gentle measurement lemma [Aar18].
@@ 10.2 Extraction of Preimages via LWE Search-Decision Reduction
$\mathop{\mathsf{Ext}}$ runs in time $T^{\prime}=O(T\cdot knB_{\mathbf{x}}\cdot\mathsf{poly}(\log(1/\delta),1/\epsilon))$, where $T$ is the running time of the decryptor $\rho_{\mathsf{Delete}}$ and $B_{\mathbf{x}}=\mathsf{max}_{\mathbf{x}_{i,j}}|\mathbf{x}_{i,j}|$ and $\mathbf{x}_{i,j}$.
@@
10.2.1 Proof Outline

### 06 Mar 2024 v4

@@ 3 Acknowledgement
We thank an anonymous reviewer who points out an issue in the analysis of the LWE search-to-decision extractor analysis in a previous version of this paper, which has now been fixed.
@@ 5.1 Projective Implementation and Threshold Implementation
Motivated by the discussion above, [Zha20] (inspired by [MW05]) formalizes a new measurement procedure for testing a state received by an adversary. We will be adopting this procedure when defining security of secure key leasing schemes in Section 8. Consider the following procedure as a binary POVM ${\mathcal{P}}$ acting on an alleged quantum-decryptor program $\rho$: sample a ciphertext $x$, evaluates the quantum decryptor on $x$, and checks if the output is correct. In a nutshell, the new procedure consists of applying an appropriate projective measurement which measures the success probability of the tested state $\rho$ under ${\mathcal{P}}$, and to output “accept” if the success probability is high enough. Of course, such measurement will not be able extract the exact success probability of $\rho$, as this is impossible from we have argued in the discussion above. Rather, the measurement will output a success probability from a finite set, such that the expected value of the output matches the true success probability of $\rho$. We will now describe this procedure in more detail.
@@ 5.2 Approximating Threshold Implementation
Our algorithm is parameterized by a distribution ${\mathcal{D}}$, collection of projective measurements ${\mathcal{P}}$, and real values $0<\epsilon,\delta,\gamma\leq 1$, and is denoted as ${\sf ATI}_{{\mathcal{P}},{\mathcal{D}},\gamma}^{\epsilon,\delta}$. The running time of the following algorithm in Figure 1 is $\frac{2\ln(4/\delta)}{\epsilon^{2}}$.
Figure 1: ATI Algorithm
@@ 5.2.1 Identity-like Approximate Threshold Implementation
  • For any input state $\rho$, let $\rho^{\prime}$ be the state after applying ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},1/2+\gamma}$ on $\rho$, we have $\lVert\rho-\rho^{\prime}\rVert_{\operatorname{Tr}}\leq O(\eta\cdot\frac{\ln(4/\delta)}{\epsilon})$  [^8] In our applications, $\eta\cdot\frac{\ln(4/\delta)}{\epsilon}$ is inverse exponential in security parameter $\lambda$..
@@
We obtain this theorem by making observations on the state after running the ATI algorithm, see Figure 1.
We will use Claim 6.3 in [Zha20], and some explanations under theorem 6.2 in [Zha20].(For the clarity of presentation, we will integrate [Zha20]’s claims into our context instead of citing it directly.)
We can write the input state at the beginning as $\rho=\sum_{i}\lambda_{i}\ket{\phi_{i}}\bra{\phi_{i}}$ |, where $p_{i}$ is an eigenvalue with eigenvector |$\ket{\phi_{i}}$ for measurement ${\mathcal{P}}_{{\mathcal{D}}_{0},{\mathcal{D}}_{1}}$ in Definition 5.11(see discussions inSection 5.1 and Lemma 5.2). In other words, ${\mathcal{P}}_{{\mathcal{D}}_{0},{\mathcal{D}}_{1}}\ket{\phi_{i}}=p_{i}\ket{\phi_{i}}$. Meanwhile, {|$\{\ket{\phi_{i}}\}_{i}$ form a basis, and the density operator $\rho$ can always be written in terms of coefficients $\{\lambda_{i}\}_{i}$ with their corresponding state in {|$\{\ket{\phi_{i}}\}_{i}$ because such a spectral decomposition always exists for any $\rho$.
@@
We consider the state after doing $T$ number of iterations of the main loop in ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},1/2+\gamma}$ algorithm (see Figure 1). Let $L=(L_{1},L_{2},\cdots,L_{2T})$ be the sequence of measurement outcomes we obtain in the ATI algorithm from the alternating tests ${\sf CProj}_{{\mathcal{P}}}$ and ${\sf IsUniform}$; let $t$ be the number of $i$ such that $L_{i-1}=L_{i}$ we have in $L$.
Since at the end of the algorithm ${\sf ATI}$, we will stop at a ${\sf IsUniform}$ test with measurement outcome 1, we obtain the following state state, as claimed by [Zha20] Claim 6.3.
$\rho^{\prime}=\sum_{i}\lambda_{i}\cdot\frac{p_{i}^{t}\cdot(1-p_{i})^{2T-t}}{\sum_{i}\lambda_{i}\cdot p_{i}^{t}\cdot(1-p_{i})^{2T-t}}\ket{\phi_{i}}\bra{\phi_{i}}$ | (3)
[Zha20] gave the form of an unnormalized pure state w.l.o.g.; we can easily derive that the form of a (normalized) mixed state evolved from its initilized form $\rho=\sum_{i}\lambda_{i}\ket{\phi_{i}}\bra{\phi_{i}}$ | will be the above.
@@
Here $\tilde{p}=\frac{t}{2T}$ is the value we obtain at the next-to-last step in ATI algorithm in Figure 1 (before we check $\tilde{p}\geq 1/2+\gamma$); $p$ is an outcome we obtain if we apply an ideal projective implementation of ${\mathcal{P}}$– in short we obtain some $p=p_{i^{*}}$ with probability $\sum_{j}\lambda_{j}$ where all |$\ket{\phi_{j}}$ have eigenvalues $p_{j}=p_{i^{*}}$ under ${\mathcal{P}}$ (see discussions in Section 5.1). But we don’t need to worry about the detailed distribution of $p$ here: since we must have $p\in[1/2,1/2+\eta]$, we can deduce that $\frac{(1-\epsilon)}{2}\cdot 2T\leq t\leq(\frac{(1+\epsilon)}{2}+\eta)\cdot 2T$. Since we have $\eta<\epsilon/2$ by our assumption and let $T=\ln(4/\delta)/\epsilon^{2}$ we have $T-\frac{\ln(4/\delta)}{\epsilon}\leq t\leq T+\frac{2\ln(4/\delta)}{\epsilon}$, with probability $1-\delta$. Now we analyze the trace distance between $\rho,\rho^{\prime}$. Without loss of generality, first assume $t\geq T$ and $\lambda_{i}-\lambda_{i}\cdot\frac{p_{i}^{t}\cdot(1-p_{i})^{2T-t}}{\sum_{i}\lambda_{i}\cdot p_{i}^{t}\cdot(1-p_{i})^{2T-t}}\geq 0,\forall i$.
@@
$\leq\frac{1}{2}\sum_{i}\lambda_{i}(1-(1-4\frac{\ln(4/\delta)}{\epsilon^{2}}\cdot\eta^{2}-8\eta\cdot\frac{2\ln(4/\delta)}{\epsilon}+4\eta^{2}\cdot 8\eta\cdot\frac{2(\ln(4/\delta))^{2}}{(\epsilon^{3})}))$
@@
The second to last inequality is because $T=\frac{\ln(4/\delta)}{\epsilon^{2}}$, and $t-T\leq\frac{\ln(4/\delta)}{\epsilon}$.
@@
$\lVert\rho-\rho^{\prime}\rVert_{\operatorname{Tr}}$ $\leq\frac{1}{2}\sum_{i}\lambda_{i}((1-4\eta^{2})^{-T}\cdot(\frac{1/2+\eta}{1/2-\eta})^{2(t-T)}-1)$
$\leq\frac{1}{2}\sum_{i}\lambda_{i}((1-4T\eta^{2})^{-1}\cdot(1-4\eta(2t-2T))^{-1}-1)$
$\leq\frac{1}{2}\frac{O(\eta\cdot\frac{\ln(4/\delta)}{\epsilon})}{1-O(\eta\cdot\frac{\ln(4/\delta)}{\epsilon})}$
@@
where the last inequality is because $\eta<\epsilon$ and $\ln(4/\delta)$ is inverse polynomial; therefore there exists some constant $10<c<$ 1 such that $1-c\leq 1-O(\eta\cdot\frac{\ln(4/\delta)}{\epsilon})$.
@@
Suppose the statistical distance of ${\mathcal{D}}_{0},{\mathcal{D}}_{1}$ is 0. ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},1/2+\gamma}$ is the ATI for mixture of projections ${\mathcal{P}}=({\mathcal{P}}_{{\mathcal{D}}_{0},{\mathcal{D}}_{1}},{\mathcal{Q}}_{{\mathcal{D}}_{0},{\mathcal{D}}_{1}})$ defined in Definition 5.11, with parameters $\gamma,\epsilon,\delta$ where $\gamma$ is inverse polynomial and $\epsilon<\gamma$. For any input state $\rho$, let $\rho^{\prime}$ be the state after applying ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},1/2+\gamma}$ on $\rho$, with probability $(1-\delta)$, we have that $\lVert\rho-\rho^{\prime}\rVert_{\operatorname{Tr}}=0$ and ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},1/2+\gamma}$ outputs 0.
@@ 7.2.4 Efficient Range Preparation for NTCF with Small LWE Secrets
Next, we measure the last register to obtain $\mathbf{y}=\mathbf{x}_{b}\mathbf{A}+\mathbf{e}^{\prime}+\mathbf{b}\cdot\mathbf{s}\mathbf{A}$. We can then uncompute the register containing $\mathbf{e}^{\prime}$ using the information in register containing $b,\mathbf{x}$, the key $(\mathbf{A},\mathbf{u})$ and the last register.
@@ 8.2 Strong SKL-PKE PKE Security: Threshold Implementation Version
In this section, we define a security notion we call Strong SKL-PKE, which is described via the measurement ${\sf TI}$ in Section 5. We show that it implies the regular security notion in the previous section. Note that Strong SKL-PKE is the security definition we prove our scheme’s security with respect to.
@@ 9.1 Parameters
    The ratios $\frac{B_{V}}{B_{L}}$, $\frac{B_{P}}{B_{V}}$, $\frac{B_{P^{\prime}}}{B_{P}}$, $\frac{B_{X}}{B_{S}}$ are super-polynomial in $\lambda$.
@@ 9.3 Correctness
Note that we can write out the entire quantum key as ${\sf\rho_{\mathsf{sk}}}=\\ \frac{1}{\sqrt{2^{k}}}\sum_{\mathbf{b}_{j}\in\{0,1\}^{k}}\lvert\mathbf{b}_{j,1},\mathbf{x}_{k,\mathbf{b}_{j,1}},\cdots,\mathbf{b}_{j,k},\mathbf{x}_{1,\mathbf{b}_{j,k}}\rangle$
By applying the decryption procedure, we will have:
@@
Since we have $\lVert\sum_{i}(\mathbf{e}_{i}^{\prime}+\mathbf{e}_{i})\rVert\leq 2k\cdot\sqrt{m}\cdot B_{P},\lVert\mathbf{e}^{\prime\prime}\rVert\leq\sqrt{m}\cdot B_{P^{\prime}}$ and $\lVert(\sum_{i\in[k]}(-\mathbf{b}_{j,i}\cdot\mathbf{e}_{i}+\mathbf{e}_{i}^{\prime})\cdot\mathbf{R}+\mathbf{e}^{\prime\prime})\mathbf{G}^{-1}(\mathbf{0}|\lfloor\frac{q}{2}\rfloor)\rVert\leq\frac{q}{4}$ for all support $\mathbf{b}_{j}\in\{0,1\}^{k}$ in the final outcome, we will obtain $\mu$ with probability $(1-O(e^{-q/B_{P^{\prime}}}))$ for all support in the above state. We can write the final output for $\mu$ in the third (rightmost) register and use $\mathsf{ct}$ to uncompute the second register and recover ${\sf\rho_{\mathsf{sk}}}$.
@@
Reusability of the decryption key follows from correctness with probability $(1-O(e^{-q/B_{P^{\prime}}}))$ where the exponent itself $-\frac{q}{B_{P^{\prime}}}$ is subexponentially (and of course superpolynomially) small. Therefore, we can apply the gentle measurement lemma [Aar18] after each honest decryption, for arbitrarily polynomial times and the error still remains negligible.
@@ 10.2 Extraction of Preimages via LWE Search-Decision Reduction
$\mathop{\mathsf{Ext}}$ runs in time $T^{\prime}=O(T\cdot knB_{\mathbf{x}}\cdot\mathsf{poly}(\log(1/\delta),1/\epsilon))$, where $T$ is the running time of the decryptor $\rho_{\mathsf{Delete}}$ and $B_{\mathbf{x}}$ is the maximum value for each entry $\mathbf{x}_{i,j}$ in the secret $\mathbf{x}$.
@@
10.2.1 Proof Outline Using Inefficient Measurement ${\sf TI}$